OpenStudy (anonymous):

A steel reinforcing rod is 12 ft. long and 1" thick in diameter. If a cubic foot of steel weighs 490 lb., how much does the rod weigh (to the nearest pound)?

OpenStudy (yanasidlinskiy):

Clues: Given dimensions of a rod... perhaps could calculate something about the rod weight of a cubic foot of steel: "Cubic foot"... sounds like a volume question!

OpenStudy (yanasidlinskiy):

How much steel is in that rod? Can you find its volume, which would be measured in cubic feet? If you knew how many cubic feet of steel is in the rod, and you know how much a single cubic foot of steel weighs, then you can multiply and get the total weight.

OpenStudy (yanasidlinskiy):

Oops.. It's 1 inch not 1 foot. like I marked it.

OpenStudy (anonymous):

The formula to find the volume is V = (pi)r^2*h ... but the answer I'm getting isn't working :P

OpenStudy (yanasidlinskiy):



OpenStudy (anonymous):

so... .5/12 squared?

OpenStudy (yanasidlinskiy):

No. It is 490 pounds. That's for 1 cubic foot, but you need to find out how much volume is in the rod... it could be more than or less than 1 cubic foot. volume of cylinder = (area of the circle that is the base & top) X height But you have to use consistent units... can't mix inches and feet. Need to convert radius to feet.

OpenStudy (phi):

**** so... .5/12 squared? ***** if you mean that is r^2, yes. You can also write it \[r= \frac{1}{2} \cdot \frac{1}{12} = \frac{1}{24} \text{ ft} \] now use that in the equation for the volume \[ V = \pi r^2 h \\ = \pi \left(\frac{1}{24}\right)^2 \cdot 12 \]

OpenStudy (phi):

the answer will be in cubic feet. to find the weight, use Volume (in cubic ft) *490 pounds per cubic ft

OpenStudy (yanasidlinskiy):

So am I right or wrong? phi..

OpenStudy (anonymous):

So is it 10 lbs?

OpenStudy (phi):

what did you get for the volume ?